OpenStudy (anonymous):

according to faraday's law voltage can be changed by changing the magnetic field strength true or false?

OpenStudy (anonymous):

\[e.m.f.=-\frac{ d \Phi }{ dt }=-\frac{ d (\bar B·\bar S) }{ dt }=-\bar S·\frac{ d \bar B }{ dt }-\bar B·\frac{ d \bar S }{ dt }\] B stands for Magnetic Field , therefore both its strength and its variations versus time will impact induced electromotive forces. I would say TRUE

OpenStudy (anonymous):

awsome thank you! :)